The stopcock connecting a 3.86 L bulb containing oxygen gas at a pressure of 6.50 atm, and a 6.74 L bulb containing helium gas at a pressure of 4.26 atm, is opened and the gases are allowed to mix. Assuming that the temperature remains constant, the final pressure in the system is atm Submit Answer Retry Entire Group 9 more group attempts remaining
Answer #1

PV1 = 3.86*6.5

PV2 = 6.74*4.26

PVtotal = PV1 + PV2 = 25.09+28.7124 = 53.8024

if V final = 3.86 + 6.74 = 10.6

PVtotal = 53.8024

P*10.6 = 53.8024

P = 53.8024/10.6

P = 5.0756 atm

which makes sense, since it is between Pressures given